OpenStudy (anonymous):

|3x + 12| = 18 SOLVE

OpenStudy (here_to_help15):

What times 3 can equal 18 ?

OpenStudy (anonymous):

6

OpenStudy (here_to_help15):

Correct so x=6

OpenStudy (here_to_help15):

Wait wait wait

OpenStudy (anonymous):

Thats not even close to one of the options

OpenStudy (here_to_help15):

I made a huge mistake lol

OpenStudy (rock_mit182):

what are you talking about

OpenStudy (rock_mit182):

you have to solve this : 3x + 12 = 18 and 3x + 12 =-18

OpenStudy (rock_mit182):

because is an absolute value

OpenStudy (anonymous):

2 is one of them right

OpenStudy (rock_mit182):

now solve the second equation

OpenStudy (anonymous):

-10

OpenStudy (here_to_help15):

Ok so I'll break it down 3x+12= 18 -12 -12 -------------- 3x=6 \[\frac{ 6 }{ 3 } =?\]

OpenStudy (rock_mit182):

now you're right, that means how far are you from certain point

OpenStudy (here_to_help15):

This is an equation @rock_mit182 the left side is suppose to equal the right side....

OpenStudy (rock_mit182):

and you figure this out by taking what is inside | | = a and what is inside of | | = -a

OpenStudy (rock_mit182):

@Here_to_help15 here we are talking about absolute valu functions

OpenStudy (rock_mit182):

if we have |x| = 3

OpenStudy (rock_mit182):

that means how far you're from zero

OpenStudy (here_to_help15):

x does not equal 3!

OpenStudy (rock_mit182):

which could be -3 or 3 units

OpenStudy (rock_mit182):

Im not talking about the last example

OpenStudy (rock_mit182):

by the way, Im just explaining to you what means an absolute value

OpenStudy (here_to_help15):

We are both correct....

OpenStudy (here_to_help15):

x=2 or -10

OpenStudy (rock_mit182):

of course we do.. exactly

OpenStudy (rock_mit182):

(y)

OpenStudy (here_to_help15):

So @brooke_ivy99 which answer choice do you have?